OpenStudy (anonymous):

Does x^4+9=90 x=3? Thanks. Does x^4+9=90 x=3? Thanks. @Mathematics

OpenStudy (anonymous):

x^4 = 90-9 x^4 = 81 x = + 3, -3

OpenStudy (anonymous):

Alright thanks! How about this one? (x-3)^5=60 Is it x=5.27?

myininaya (myininaya):

3i and -3i are also solutions to that first one

myininaya (myininaya):

\[x-3=60^\frac{1}{5} => x=3+60^\frac{1}{5}\]

myininaya (myininaya):

which is approximately what you have

OpenStudy (anonymous):

Thank you, is that the only answer?

myininaya (myininaya):

you can write it in polar form to find other solutions if they exist